OpenStudy (anonymous):

For equilateral triangle: Find x if line segment XZ = 1/2x + 8 and ZY = 2(x-8)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

set them equal to each other 1/2x +8 = 2x -16 -3/2x = -24 x = 16

OpenStudy (anonymous):

I am assuming that the triangle is labeled XYZ. Remember that an equilateral triangle has three congruent (or equal) sides. Thus, side XZ = side ZY. Set the measures of the two sides equal to each other and solve for x. 0.5x + 8 = 2(x-8) 0.5x = 2x - 16 - 8 -1.5x = -24 x = 16